### v1

@@ 3.1 Validation
Table 1: Metric-guided bounds yield valid coverages while conventional pixel-wise bounds do not. Using 20 patients and target coverage of 90%, we perform leave-one-out cross-validation and compute average coverage using metric-guided and pixel-wise methods for maximum dose to the heart (Heart $D_{0}$), volume of ipsilateral lung that received 20Gy (Right Lung $V_{20}$), volume of ipsilateral lung (Right Lung Volume), and dose that 35% volume of the ipsilateral lung receives (Right Lung $D_{35}$).
Method Heart $D_{0}$ Right Lung Volume Right Lung $V_{20}$ Right Lung $D_{35}$
Metric-guided 90 90 90 90
Pixel-wise 75 0 50 50
3.2 Upper and lower bound retrieval
We retrieve metric-guided and pixel-wise upper and lower bounds for a target coverage of 90% for maximum dose to the heart (Fig. 2). To verify the retrieved images are representative of the bounds at test time, we compute retrieval error defined as:
@@ 3.2 Upper and lower bound retrieval
where $B$ denotes the calibrated bound and can be upper bound $UB$ or lower bound $LB$, and $\hat{Y}_{B}^{n_{p}+1}=\operatorname*{arg\,min}_{\hat{Y}_{j}^{n_{p}+1}}{|\hat{Y}_{j}^{n_{p}+1}-B^{n_{p}+1}|}$ are the estimated metrics closest to the calibrated bounds. We find that pixel-wise upper and lower bounds are perceptually similar and only differ in their intensity, while metric-guided bounds differ in the spatial distribution of pixel intensities. This indicates that metric-guided bounds take spatial correlations into account. As a consequence, the pixel-wise differences for metric-guided bounds can be both positive and negative. This indicates that single pixels do not carry sufficient information to explain the variations in dose. We find that the segmentations of the heart are also perceptually different. Pixel-wise upper bounds tend to have larger volumes than lower bounds, while this rule does not hold for metric-guided bounds.
Furthermore, this result suggests that pixel-wise and metric-guided methods may disagree on inliers and outliers. Metric-guided inlier reconstructions may have pixels considered as pixel-wise outliers and metric-guided outlier reconstructions may have pixels considered as pixel-wise inliers.
3.3 Anatomical Differences
@@
Using organ segmentations from RPA, we determine whether there is a statistically significant difference in upper bound volume across methods and lower bound volume across methods using paired t-tests. We use the dose metrics in Table 1. We find statistically significant differences (p<0.05) for upper and lower bounds across methods except for the upper bound reconstructions for Heart $D_{0}$ (p=7.2e-2) (Fig. 3). This suggests that the upper and lower bounds across methods are anatomically different.
@@
4 Conclusion
We propose a method that leverages conformal prediction to retrieve upper/lower bounds and statistical inliers/outliers of reconstructions based on the prediction intervals of downstream metrics. We apply our method to sv-CT for downstream radiotherapy planning and show 1) metric-guided bounds have valid coverage for downstream metrics unlike conventional pixel-wise bounds and 2) statistically significant anatomical differences of upper/lower bounds between metric-guided and pixel-wise methods.
5 Discussion
There are several areas for further investigation:
Factors affecting retrieval error. Retrieval error may depend on number of samples, the diversity of samples, and the accuracy of the model. The prediction intervals and retrieval errors may also be very large if the model is highly biased. Asymmetric bounds could help identify this bias [26]. Furthermore, we assume the downstream processes to be deterministic. This is an appropriate assumption for the maximum dose to the heart, but may not be for other parameters. Opportunities lie in decoupling uncertainty from physical, reconstruction algorithm, and downstream algorithm randomness [13].
Evaluating Safety and Equity. We can perform patient-specific safety evaluations and identify inequities across patients. For a dose prescription of 50Gy (2Gy/fraction), a safe maximum dose to the heart is <5Gy and the volume of the ipsilateral lung getting 20Gy is <35%. If the upper bound of the prediction interval is greater than these thresholds, it may indicate that the reconstruction is unsuitable for planning. Patients or measurement conditions with high uncertainty can be used for downstream interpretation [25, 22] and action [34, 20]. They may correspond to specific clinical scenarios, such as inadequately filled lungs or large distance from heart to chest wall. Opportunities lie in applying causal methods [24, 28, 23] to identify factors causes of high uncertainty.
Test time evaluation metrics for reconstruction. While we show inliers and outliers for one metric, our method can be extended to multiple metrics [21, 31] where we find reconstructions with all estimated metrics in the prediction intervals containing the ground truth metrics with confidence. Opportunities lie in assessing reconstructions with multiple critical metrics.
Other applications. Opportunities lie in extending our method to other medical imaging applications [22, 17] and critical scenarios. Additionally, although not demonstrated in our work, our method does not necessitate reconstruction samples to be of identical size or dimensions, as calibration is conducted based on a scalar downstream metric.
{credits}
5.0.1 Acknowledgements

### v2

@@ 4 Results
Table 1: Metric-guided bounds yield valid coverages while conventional pixel-wise bounds do not. Using 20 patients and target coverage of 90%, we perform leave-one-out cross-validation and compute average test coverage, interval length, and probability of reconstruction classified as an outlier based on prediction intervals (P(outlier)) using metric-guided (symmetric and asymmetric adjustments) and pixel-wise methods for maximum dose to the heart (Heart $D_{0}$), heart volume, volume of ipsilateral lung that received 20Gy (Right Lung $V_{20}$), dose that 35% volume of the ipsilateral lung receives (Right Lung $D_{35}$), and volume of ipsilateral lung (Right Lung Volume).
4.1 Interpreting Pixel-wise and Metric-guided Bounds
@@
We show that the upper/lower bounds can be interpretated (Fig. 2). We find that pixel-wise upper and lower bounds are perceptually similar and only differ in their intensity, while metric-guided bounds differ in the spatial distribution of pixel intensities. This indicates that metric-guided bounds take spatial correlations into account. As a consequence, the pixel-wise differences for metric-guided bounds can be both positive and negative. This indicates that single pixels do not carry sufficient information to explain the variations in dose. We find that the segmentations of the heart are also perceptually different. Pixel-wise upper bounds tend to have larger volumes than lower bounds, while this rule does not hold for metric-guided bounds. Furthermore, this result suggests that pixel-wise and metric-guided methods may disagree on inliers and outliers. Metric-guided inlier reconstructions may have pixels considered as pixel-wise outliers and metric-guided outlier reconstructions may have pixels considered as pixel-wise inliers.
Figure 2: Metric-guided bounds account for spatial correlations that affect downstream metrics. For maximum dose to the heart (D) with target coverage of 90%, we show contours for heart (red), right lung (blue), left lung (yellow), and body (green) overlaid on CT slices. Pixel-wise upper and lower bounds differ in pixel-wise intensity, while metric-guided bounds differ in the spatial distribution of pixel intensities. Pixel-wise upper bounds have larger heart volumes than lower bounds, while metric-guided bounds have similar heart volumes. Retrieval error $\varepsilon_{B}$ is the difference between estimated and actual bound divided by the interval length (Eq. 2). We used symmetric adjustments for metric-guided bounds.
@@
Furthermore, we explore the anatomical differences between pixel-wise and metric-guided bounds (Fig. 3). Using organ segmentations from RPA, we determine whether there is a statistically significant difference in upper bound volume across methods and lower bound volume across methods using paired t-tests. We use the dose metrics in Table 1. We find statistically significant differences (p<0.05) for upper and lower bounds across methods except for the upper bound reconstructions for Heart $D_{0}$ (p=7.2e-2) (Fig. 3). This suggests that the upper and lower bounds across methods are anatomically different.
@@
5 Conclusion and Discussion
We propose a method that leverages conformal prediction to retrieve upper/lower bounds and statistical inliers/outliers of reconstructions based on the prediction intervals of downstream metrics. We apply our method to sparse CT for downstream radiotherapy planning and show 1) metric-guided bounds have valid coverage for downstream metrics unlike conventional pixel-wise bounds and 2) statistically significant anatomical differences of upper/lower bounds between metric-guided and pixel-wise methods. There are many areas for further investigation:
  1. 1.
    Factors affecting retrieval error. Retrieval error may depend on number of samples, the diversity of samples, and the accuracy of the model. We suspect that the larger the number and the diversity of samples, the better the approximation. The prediction intervals and retrieval errors may also be very large if the model is highly biased. Asymmetric bounds could help identify this bias. Furthermore, we assume the downstream processes to be deterministic. This is an appropriate assumption for the maximum dose to the heart, but may not be for other parameters. Opportunities lie in decoupling uncertainty from physical, reconstruction algorithm, and downstream algorithm randomness [37].
  2. 2.
    Evaluating Safety and Equity. We can perform patient-specific safety evaluations and identify inequities across patients. For a dose prescription of 50Gy (2Gy/fraction), a safe maximum dose to the heart is <5Gy and the volume of the ipsilateral lung getting 20Gy is <35%. If the upper bound of the prediction interval is greater than these thresholds, it may indicate that the reconstruction is unsuitable for planning. The larger the prediction interval length, the higher the uncertainty. Patients or measurement conditions with high uncertainty can be used for downstream interpretation [38] and action [39, 40]. They may correspond to specific clinical scenarios, such as inadequately filled lungs or large distance from heart to chest wall. Interestingly, we find that pixel-wise volume and intensity intervals for the right lung can explain some of the variance of metric-guided dose intervals for $V_{20}$ ($R^{2}$=0.801) and $D_{35}$ ($R^{2}$=0.241). Opportunities lie in applying causal methods [41, 42, 43] to identify factors causes of high uncertainty.
  3. 3.
    Multiple metrics. Our method is capable of considering multiple metrics and find reconstructions with all estimated metrics in the prediction intervals containing the ground truth metrics with confidence. While we use a complete set of projections as ground truth, we can also use image quality criteria to assess whether reconstructions meet the measurement standards [44, 45]. Opportunities lie in assessing reconstructions with multiple critical metrics.
  4. 4.
    Other applications. Opportunities lie in extending our method to other medical imaging applications [46] and critical scenarios. Additionally, although not demonstrated in our work, our method does not necessitate reconstruction samples to be of identical size or dimensions, as calibration is conducted based on a scalar downstream metric.
Acknowledgements